# The theory of implicit operations

### v1

@@ 12. Structure theory
In this section, we address the following question: what do we gain by moving from a quasivariety to its Beth companion? We will answer it by showing that not only does the Beth companion of a quasivariety $\mathsf{K}$ inherit a significant portion of the structure theory of $\mathsf{K}$, but its structure theory is often much richer than that of $\mathsf{K}$. To this end, it is convenient to restrict our attention to the pp expansions that respect the structure of congruence lattices, which we term congruence preserving. More precisely, recall that if $\mathsf{M}$ is a pp expansion of a quasivariety $\mathsf{K}$, then $\mathsf{M}$ is a quasivariety and $\mathsf{K}$ the class of ${L}_{\mathsf{K}}$-subreducts of $\mathsf{M}$ (see 10.3(ii) and 10.2).
@@
A pp expansion $\mathsf{M}$ of a class of algebras $\mathsf{K}$ is said to be equational when it is faithfully term equivalent relative to $\mathsf{K}$ to a pp expansion of $\mathsf{K}$ induced by some $\mathcal{F}\subseteq\mathsf{ext}_{eq}(\mathsf{K})$ and ${L}_{\mathcal{F}}$.
@@
We say that a Beth companion is congruence preserving when it is a congruence preserving pp expansion. Similarly, we call a Beth companion equational when it is an equational pp expansion. It is an immediate consequence of 11.12(vi) that a Beth companion of a class of algebras $\mathsf{K}$ is equational if and only if it is faithfully term equivalent relative to $\mathsf{K}$ to a Beth companion of $\mathsf{K}$ induced by some $\mathcal{F}\subseteq\mathsf{ext}_{eq}(\mathsf{K})$ and ${L}_{\mathcal{F}}$. In view of 12.4, all the Beth companions considered so far are congruence preserving as we proceed to illustrate.
@@
Notice that all the Beth companions of varieties mentioned in 11.9 are equational (in fact, they are all induced by sets of implicit operations defined by conjunctions of equations). Therefore, 12.4 guarantees that they are congruence preserving. ∎
@@
For every $n\in\mathbb{Z}^{+}$ let $\bm{A}_{n}$ be the unique Heyting algebra whose lattice reduct is obtained by adding a new maximum $1$ to the powerset lattice $\langle\mathcal{P}(\{1,\dots,n\});\cap,\cup\rangle$. Then for every $n\geqslant 3$ the variety $\mathbb{V}(\bm{A}_{n})$ admits a Beth companion that is congruence preserving and not equational, as shown in [27, Thm. 3.3]. Notably, the variety $\mathbb{V}(\bm{A}_{n})$ can be equivalently described as the variety of Heyting algebras of depth $\leqslant 2$ (see 10.18) satisfying the bounded width $\leqslant n$ axiom in the variables $x_{1},\dots,x_{n+1}$, namely, the equation
@@
In addition, congruence preserving pp expansions preserve and reflect the property of “being relatively (finitely) subdirectly irreducible” (cf. 10.3(ii)) and preserve the property of “being a variety”. We recall that the latter fails for arbitrary pp expansions (see [27, Thm. 2.1]). More precisely, we will prove the following.
@@
In view of 12.10, under reasonable assumptions, the structure theory of a congruence preserving Beth companion $\mathsf{M}$ of a quasivariety $\mathsf{K}$ enhances that of $\mathsf{K}$ as follows: $\mathsf{M}$ turns out to be a variety (as opposed to a quasivariety) which, moreover, is arithmetical (as opposed to relatively congruence distributive only) and possesses the congruence extension property.
Let us illustrate the effect of 12.10 in the setting of filtral quasivarieties. A variety $\mathsf{K}$ is said to be discriminator when there exist a class of algebras $\mathsf{M}$ and a quaternary term $t$ such that $\mathsf{K}=\mathbb{V}(\mathsf{M})$ and $t^{\bm{A}}$ is the quaternary discriminator function on $A$ for every $\bm{A}\in\mathsf{M}$ [22, 107] (see also [21, Sec. IV.9]). Examples of discriminator varieties include the variety of Boolean algebras and for each $n\in\mathbb{N}$ the variety of rings satisfying the equation $x\thickapprox x^{n}$ (see, e.g., [20, pp. 179–180]). The importance of discriminator varieties derives from the fact that they admit a general representation theorem in terms of Boolean products with subdirectly irreducible factors [21, Thm. IV.9.4] (see also [104]). While every discriminator variety is filtral (see, e.g., [10, p. 101]), the converse need not hold in general: for instance, while the variety of (bounded) distributive lattices is filtral, it is not a discriminator variety. However, its Beth companion, the variety of Boolean algebras, is a discriminator variety. From 12.10 we will infer that this is true in general.
@@
As the class of integral domains of characteristic zero is closed under nontrivial subalgebras, from 12.10 it follows that $\mathsf{Meadow}_{0}$ is an arithmetical variety with the congruence extension property. None of these facts holds for $\mathsf{RCRing}_{0}$, which is a proper quasivariety without the relative congruence extension property (congruence permutability is a property of varieties only), as we proceed to illustrate.
@@
Next we prove Theorems 12.4, 12.7, 12.9, and 12.10 and Corollary 12.11. The proof of 12.4 hinges upon the next pair of observations.
@@
Let $\mathsf{M}$ be a pp expansion of a quasivariety $\mathsf{K}$ induced by $\mathcal{F}\subseteq\mathsf{ext}_{pp}(\mathsf{K})$ and ${L}_{\mathcal{F}}$. Then $\mathsf{Con}_{\mathsf{M}}(\bm{A})=\mathsf{Con}_{\mathsf{K}}(\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}})$ for every $\bm{A}\in\mathsf{K}[{L}_{\mathcal{F}}]$.
@@
Consider $\bm{A}\in\mathsf{K}[{L}_{\mathcal{F}}]$. As the inclusion $\mathsf{Con}_{\mathsf{M}}(\bm{A})\subseteq\mathsf{Con}_{\mathsf{K}}(\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}})$ always holds (see Remark 12.2), we detail only the proof of the reverse inclusion. Consider $\theta\in\mathsf{Con}_{\mathsf{K}}(\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}})$. Then $(\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}})/\theta\in\mathsf{K}$. Since $\mathcal{F}\subseteq\mathsf{ext}_{pp}(\mathsf{K})$ and $\mathsf{K}$ is a quasivariety, by 9.6 there exists $\bm{B}\in\mathsf{K}[{L}_{\mathcal{F}}]$ such that $(\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}})/\theta\leqslant\bm{B}{\upharpoonright}_{{L}_{\mathsf{K}}}$. Therefore, the canonical surjection $h\colon\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}}\to\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}}/\theta$ can be viewed as a homomorphism $h\colon\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}}\to\bm{B}{\upharpoonright}_{{L}_{\mathsf{K}}}$. As $\bm{A},\bm{B}\in\mathsf{K}[{L}_{\mathcal{F}}]$, Proposition 9.5 guarantees that $h$ is also a homomorphism from $\bm{A}$ to $\bm{B}$. Together with $\bm{B}\in\mathsf{M}$ and $\mathsf{Ker}(h)=\theta$, this yields that $\theta$ is an $\mathsf{M}$-congruence of $\bm{A}$. ∎
@@
(i): Since $\mathsf{M}$ is equational, there exists a pp expansion $\mathsf{N}$ of $\mathsf{K}$ induced by some $\mathcal{F}\subseteq\mathsf{ext}_{eq}(\mathsf{K})$ and ${L}_{\mathcal{F}}$ such that $\mathsf{M}$ is faithfully term equivalent to $\mathsf{N}$ relative to $\mathsf{K}$. We can apply 10.4, obtaining $\mathsf{N}=\mathsf{K}[{L}_{\mathcal{F}}]$. Therefore, $\mathsf{Con}_{\mathsf{N}}(\bm{A})=\mathsf{Con}_{\mathsf{K}}(\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}})$ for every $\bm{A}\in\mathsf{N}$ by 12.13. Hence, we conclude that $\mathsf{N}$ is congruence preserving. From 12.14 it follows that $\mathsf{M}$ is congruence preserving as well.
@@
Let $\mathsf{M}$ be a pp expansion of $\mathsf{K}$ of the form $\mathbb{S}(\mathsf{K}[{L}_{\mathcal{F}}])$. From the assumption that $\mathsf{K}$ has the amalgamation property and Theorem 4.3(i) it follows that every $f\in\mathcal{F}\subseteq\mathsf{ext}_{pp}(\mathsf{K})$ is interpolated by some $f_{*}\in\mathsf{imp}_{eq}(\mathsf{K})$.
@@
of $\mathsf{K}$. We will prove that $\mathcal{F}_{*}\subseteq\mathsf{ext}_{eq}(\mathsf{K})$. As $\mathcal{F}_{*}\subseteq\mathsf{imp}_{eq}(\mathsf{K})$ by definition, it suffices to show that $\mathcal{F}_{*}\subseteq\mathsf{ext}(\mathsf{K})$. Consider an $n$-ary $f_{*}\in\mathcal{F}_{*}$, $\bm{A}\in\mathsf{K}$, and $a_{1},\dots,a_{n}\in A$. Since $f\in\mathcal{F}\subseteq\mathsf{ext}(\mathsf{K})$, there exists $\bm{B}\in\mathsf{K}$ with $\bm{A}\leqslant\bm{B}$ such that $\langle a_{1},\dots,a_{n}\rangle\in\mathsf{dom}(f^{\bm{B}})$. Consequently, $\langle a_{1},\dots,a_{n}\rangle\in\mathsf{dom}(f_{*}^{\bm{B}})$ because $f_{*}$ interpolates $f$. Hence, $f_{*}$ is extendable.
Together with Theorem 10.4, the inclusion $\mathcal{F}_{*}\subseteq\mathsf{ext}_{eq}(\mathsf{K})$ implies that $\mathsf{M}_{*}=\mathsf{K}[{L}_{\mathcal{F}_{*}}]$ is a pp expansion of $\mathsf{K}$. Therefore, to prove that $\mathsf{M}$ is an equational pp expansion of $\mathsf{K}$, it only remains to show that the pp expansions $\mathsf{M}_{*}$ and $\mathsf{M}$ are faithfully term equivalent relative to $\mathsf{K}$.
@@
Since $\mathsf{M}_{*}$ is induced by $\mathcal{F}_{*}\subseteq\mathsf{ext}_{eq}(\mathsf{K})$ and ${L}_{\mathcal{F}_{*}}$, we obtain that $\mathsf{M}$ is an equational pp expansion of $\mathsf{K}$. Then 12.4 implies that $\mathsf{M}$ is congruence preserving.
It remains to prove that $\mathsf{M}$ has the amalgamation property. Consider a pair of embeddings $f\colon\bm{A}\to\bm{B}$ and $g\colon\bm{A}\to\bm{C}$ with $\bm{A},\bm{B},\bm{C}\in\mathsf{M}$. As $\bm{B},\bm{C}\in\mathsf{M}=\mathbb{S}(\mathsf{K}[{L}_{\mathcal{F}}])$, there exist $\bm{B}^{\prime},\bm{C}^{\prime}\in\mathsf{K}[{L}_{\mathcal{F}}]$ and a pair of embeddings $i_{1}\colon\bm{B}\to\bm{B}^{\prime}$ and $i_{2}\colon\bm{C}\to\bm{C}^{\prime}$. Let $f^{\prime}\colon\bm{A}\to\bm{B}^{\prime}$ and $g^{\prime}\colon\bm{A}\to\bm{C}^{\prime}$ be the embeddings defined as $f^{\prime}=i_{1}\circ f$ and $g^{\prime}=i_{2}\circ g$. Observe that $f^{\prime}$ and $g^{\prime}$ can be viewed as embeddings $f^{\prime}\colon\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}}\to\bm{B}^{\prime}{\upharpoonright}_{{L}_{\mathsf{K}}}$ and $g^{\prime}\colon\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}}\to\bm{C}^{\prime}{\upharpoonright}_{{L}_{\mathsf{K}}}$. Furthermore, from $\bm{A},\bm{B}^{\prime},\bm{C}^{\prime}\in\mathsf{M}$ and 10.2 it follows that $\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}},\bm{B}^{\prime}{\upharpoonright}_{{L}_{\mathsf{K}}},\bm{C}^{\prime}{\upharpoonright}_{{L}_{\mathsf{K}}}\in\mathsf{K}$. Therefore, the assumption that $\mathsf{K}$ has the amalgamation property guarantees the existence of some $\bm{D}\in\mathsf{K}$ and embeddings $h_{f}\colon\bm{B}^{\prime}{\upharpoonright}_{{L}_{\mathsf{K}}}\to\bm{D}$ and $h_{g}\colon\bm{C}^{\prime}{\upharpoonright}_{{L}_{\mathsf{K}}}\to\bm{D}$ such that $h_{f}\circ f^{\prime}=h_{g}\circ g^{\prime}$.
@@
Recall that $\bm{D}\in\mathsf{K}$ and that $\mathsf{M}=\mathbb{S}(\mathsf{K}[{L}_{\mathcal{F}}])$ is a pp expansion of $\mathsf{K}$. Therefore, by 10.2 there exists $\bm{D}^{\prime}\in\mathsf{K}[{L}_{\mathcal{F}}]$ such that $\bm{D}\leqslant\bm{D}^{\prime}{\upharpoonright}_{{L}_{\mathsf{K}}}$. Let $h_{f}^{\prime}\colon\bm{B}^{\prime}{\upharpoonright}_{{L}_{\mathsf{K}}}\to\bm{D}^{\prime}{\upharpoonright}_{{L}_{\mathsf{K}}}$ and $h_{g}^{\prime}\colon\bm{C}^{\prime}{\upharpoonright}_{{L}_{\mathsf{K}}}\to\bm{D}^{\prime}{\upharpoonright}_{{L}_{\mathsf{K}}}$ be the embeddings obtained by composing $h_{f}$ and $h_{g}$, respectively, with the inclusion map of $\bm{D}$ into $\bm{D}^{\prime}{\upharpoonright}_{{L}_{\mathsf{K}}}$. Since $\bm{B}^{\prime},\bm{C}^{\prime},\bm{D}^{\prime}\in\mathsf{K}[{L}_{\mathcal{F}}]$, Proposition 9.5 implies that $h_{f}^{\prime}\colon\bm{B}^{\prime}\to\bm{D}^{\prime}$ and $h_{g}^{\prime}\colon\bm{C}^{\prime}\to\bm{D}^{\prime}$ are also embeddings. Moreover, from $h_{f}\circ f^{\prime}=h_{g}\circ g^{\prime}$ it follows that $h_{f}^{\prime}\circ f^{\prime}=h_{g}^{\prime}\circ g^{\prime}$. Together with $f^{\prime}=i_{1}\circ f$, and $g^{\prime}=i_{2}\circ g$, this yields
@@
To this end, consider $\bm{A}\in\mathsf{M}$ and observe that $\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}}\in\mathsf{K}$ by 10.2. From 2.10 it follows that
@@
It only remains to prove the last part of the statement. Suppose that $\mathsf{K}$ is a variety. From Theorem 10.3(ii) it follows that $\mathsf{M}$ is a quasivariety. Therefore, it remains to show that $\mathsf{M}$ is closed under homomorphic images. By 2.5 it suffices to prove that $\mathsf{Con}(\bm{A})=\mathsf{Con}_{\mathsf{M}}(\bm{A})$ for every $\bm{A}\in\mathsf{M}$. To this end, observe that for every $\bm{A}\in\mathsf{M}$,
@@
where the first step holds by 12.2, the second follows from $\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}}\in\mathsf{K}$ (see 10.2) and the assumption that $\mathsf{K}$ is a variety, and the third from $\bm{A}\in\mathsf{M}$ and the assumption that $\mathsf{M}$ is a congruence preserving pp expansion of $\mathsf{K}$. Thus, $\mathsf{Con}(\bm{A})=\mathsf{Con}_{\mathsf{M}}(\bm{A})$, as desired. ∎

### v2

@@ 12. Structure theory
In this section, we address the following question: what do we gain by moving from a quasivariety to its Beth companion? We will answer it by showing that not only does the Beth companion of a quasivariety $\mathsf{K}$ inherit a significant portion of the structure theory of $\mathsf{K}$, but its structure theory is often much richer than that of $\mathsf{K}$. To this end, it is convenient to restrict our attention to the pp expansions that respect the structure of congruence lattices, which we term congruence preserving. More precisely, recall that if $\mathsf{M}$ is a pp expansion of a quasivariety $\mathsf{K}$, then $\mathsf{M}$ is a quasivariety and $\mathsf{K}$ the class of ${L}_{\mathsf{K}}$-subreducts of $\mathsf{M}$ (see Theorem 10.3(ii) and Proposition 10.2).
@@
A pp expansion $\mathsf{M}$ of a class of algebras $\mathsf{K}$ is said to be equational when it is faithfully term equivalent relative to $\mathsf{K}$ to a pp expansion of $\mathsf{K}$ induced by some $\mathcal{F}\subseteq\mathsf{ext}_{\textsc{eq}}(\mathsf{K})$ and ${L}_{\mathcal{F}}$.
@@
We say that a Beth companion is congruence preserving when it is a congruence preserving pp expansion. Similarly, we call a Beth companion equational when it is an equational pp expansion. It is an immediate consequence of Remark 11.12(vi) that a Beth companion of a class of algebras $\mathsf{K}$ is equational if and only if it is faithfully term equivalent relative to $\mathsf{K}$ to a Beth companion of $\mathsf{K}$ induced by some $\mathcal{F}\subseteq\mathsf{ext}_{\textsc{eq}}(\mathsf{K})$ and ${L}_{\mathcal{F}}$. In view of Theorem 12.4, all the Beth companions considered so far are congruence preserving as we proceed to illustrate.
@@
Notice that all the Beth companions of varieties mentioned in Theorem 11.9 are equational (in fact, they are all induced by sets of implicit operations defined by conjunctions of equations). Therefore, Theorem 12.4 guarantees that they are congruence preserving. ∎
@@
For every $n\in\mathbb{Z}^{+}$ let $\bm{A}_{n}$ be the unique Heyting algebra whose lattice reduct is obtained by adding a new maximum $1$ to the powerset lattice $\langle\mathcal{P}(\{1,\dots,n\});\cap,\cup\rangle$. Then for every $n\geqslant 3$ the variety $\mathbb{V}(\bm{A}_{n})$ admits a Beth companion that is congruence preserving and not equational, as shown in [27, Thm. 3.3]. Notably, the variety $\mathbb{V}(\bm{A}_{n})$ can be equivalently described as the variety of Heyting algebras of depth $\leqslant 2$ (see Example 10.18) satisfying the bounded width $\leqslant n$ axiom in the variables $x_{1},\dots,x_{n+1}$, namely, the equation
@@
In addition, congruence preserving pp expansions preserve and reflect the property of “being relatively (finitely) subdirectly irreducible” (cf. Theorem 10.3(ii)) and preserve the property of “being a variety”. We recall that the latter fails for arbitrary pp expansions (see [27, Thm. 2.1]). More precisely, we will prove the following.
@@
In view of Theorem 12.10, under reasonable assumptions, the structure theory of a congruence preserving Beth companion $\mathsf{M}$ of a quasivariety $\mathsf{K}$ enhances that of $\mathsf{K}$ as follows: $\mathsf{M}$ turns out to be a variety (as opposed to a quasivariety) which, moreover, is arithmetical (as opposed to relatively congruence distributive only) and possesses the congruence extension property.
Let us illustrate the effect of Theorem 12.10 in the setting of filtral quasivarieties. A variety $\mathsf{K}$ is said to be discriminator when there exist a class of algebras $\mathsf{M}$ and a quaternary term $t$ such that $\mathsf{K}=\mathbb{V}(\mathsf{M})$ and $t^{\bm{A}}$ is the quaternary discriminator function on $A$ for every $\bm{A}\in\mathsf{M}$ [22, 107] (see also [21, Sec. IV.9]). Examples of discriminator varieties include the variety of Boolean algebras and for each $n\in\mathbb{N}$ the variety of rings satisfying the equation $x\thickapprox x^{n}$ (see, e.g., [20, pp. 179–180]). The importance of discriminator varieties derives from the fact that they admit a general representation theorem in terms of Boolean products with subdirectly irreducible factors [21, Thm. IV.9.4] (see also [104]). While every discriminator variety is filtral (see, e.g., [10, p. 101]), the converse need not hold in general: for instance, while the variety of (bounded) distributive lattices is filtral, it is not a discriminator variety. However, its Beth companion, the variety of Boolean algebras, is a discriminator variety. From Theorem 12.10 we will infer that this is true in general.
@@
As the class of integral domains of characteristic zero is closed under nontrivial subalgebras, from Theorem 12.10 it follows that $\mathsf{Meadow}_{0}$ is an arithmetical variety with the congruence extension property. None of these facts holds for $\mathsf{RCRing}_{0}$, which is a proper quasivariety without the relative congruence extension property (congruence permutability is a property of varieties only), as we proceed to illustrate.
@@
Next we prove Theorems 12.4, 12.7, 12.9, and 12.10 and Corollary 12.11. The proof of Theorem 12.4 hinges upon the next pair of observations.
@@
Let $\mathsf{M}$ be a pp expansion of a quasivariety $\mathsf{K}$ induced by $\mathcal{F}\subseteq\mathsf{ext}_{\textsc{pp}}(\mathsf{K})$ and ${L}_{\mathcal{F}}$. Then $\mathsf{Con}_{\mathsf{M}}(\bm{A})=\mathsf{Con}_{\mathsf{K}}(\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}})$ for every $\bm{A}\in\mathsf{K}[{L}_{\mathcal{F}}]$.
@@
Consider $\bm{A}\in\mathsf{K}[{L}_{\mathcal{F}}]$. As the inclusion $\mathsf{Con}_{\mathsf{M}}(\bm{A})\subseteq\mathsf{Con}_{\mathsf{K}}(\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}})$ always holds (see Remark 12.2), we detail only the proof of the reverse inclusion. Consider $\theta\in\mathsf{Con}_{\mathsf{K}}(\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}})$. Then $(\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}})/\theta\in\mathsf{K}$. Since $\mathcal{F}\subseteq\mathsf{ext}_{\textsc{pp}}(\mathsf{K})$ and $\mathsf{K}$ is a quasivariety, by Proposition 9.6 there exists $\bm{B}\in\mathsf{K}[{L}_{\mathcal{F}}]$ such that $(\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}})/\theta\leqslant\bm{B}{\upharpoonright}_{{L}_{\mathsf{K}}}$. Therefore, the canonical surjection $h\colon\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}}\to\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}}/\theta$ can be viewed as a homomorphism $h\colon\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}}\to\bm{B}{\upharpoonright}_{{L}_{\mathsf{K}}}$. As $\bm{A},\bm{B}\in\mathsf{K}[{L}_{\mathcal{F}}]$, Proposition 9.5 guarantees that $h$ is also a homomorphism from $\bm{A}$ to $\bm{B}$. Together with $\bm{B}\in\mathsf{M}$ and $\mathsf{Ker}(h)=\theta$, this yields that $\theta$ is an $\mathsf{M}$-congruence of $\bm{A}$. ∎
@@
(i): Since $\mathsf{M}$ is equational, there exists a pp expansion $\mathsf{N}$ of $\mathsf{K}$ induced by some $\mathcal{F}\subseteq\mathsf{ext}_{\textsc{eq}}(\mathsf{K})$ and ${L}_{\mathcal{F}}$ such that $\mathsf{M}$ is faithfully term equivalent to $\mathsf{N}$ relative to $\mathsf{K}$. We can apply Theorem 10.4, obtaining $\mathsf{N}=\mathsf{K}[{L}_{\mathcal{F}}]$. Therefore, $\mathsf{Con}_{\mathsf{N}}(\bm{A})=\mathsf{Con}_{\mathsf{K}}(\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}})$ for every $\bm{A}\in\mathsf{N}$ by Proposition 12.13. Hence, we conclude that $\mathsf{N}$ is congruence preserving. From Proposition 12.14 it follows that $\mathsf{M}$ is congruence preserving as well.
@@
Let $\mathsf{M}$ be a pp expansion of $\mathsf{K}$ of the form $\mathbb{S}(\mathsf{K}[{L}_{\mathcal{F}}])$. From the assumption that $\mathsf{K}$ has the amalgamation property and Theorem 4.3(i) it follows that every $f\in\mathcal{F}\subseteq\mathsf{ext}_{\textsc{pp}}(\mathsf{K})$ is interpolated by some $f_{*}\in\mathsf{imp}_{\textsc{eq}}(\mathsf{K})$.
@@
of $\mathsf{K}$. We will prove that $\mathcal{F}_{*}\subseteq\mathsf{ext}_{\textsc{eq}}(\mathsf{K})$. As $\mathcal{F}_{*}\subseteq\mathsf{imp}_{\textsc{eq}}(\mathsf{K})$ by definition, it suffices to show that $\mathcal{F}_{*}\subseteq\mathsf{ext}(\mathsf{K})$. Consider an $n$-ary $f_{*}\in\mathcal{F}_{*}$, $\bm{A}\in\mathsf{K}$, and $a_{1},\dots,a_{n}\in A$. Since $f\in\mathcal{F}\subseteq\mathsf{ext}(\mathsf{K})$, there exists $\bm{B}\in\mathsf{K}$ with $\bm{A}\leqslant\bm{B}$ such that $\langle a_{1},\dots,a_{n}\rangle\in\mathsf{dom}(f^{\bm{B}})$. Consequently, $\langle a_{1},\dots,a_{n}\rangle\in\mathsf{dom}(f_{*}^{\bm{B}})$ because $f_{*}$ interpolates $f$. Hence, $f_{*}$ is extendable.
Together with Theorem 10.4, the inclusion $\mathcal{F}_{*}\subseteq\mathsf{ext}_{\textsc{eq}}(\mathsf{K})$ implies that $\mathsf{M}_{*}=\mathsf{K}[{L}_{\mathcal{F}_{*}}]$ is a pp expansion of $\mathsf{K}$. Therefore, to prove that $\mathsf{M}$ is an equational pp expansion of $\mathsf{K}$, it only remains to show that the pp expansions $\mathsf{M}_{*}$ and $\mathsf{M}$ are faithfully term equivalent relative to $\mathsf{K}$.
@@
Since $\mathsf{M}_{*}$ is induced by $\mathcal{F}_{*}\subseteq\mathsf{ext}_{\textsc{eq}}(\mathsf{K})$ and ${L}_{\mathcal{F}_{*}}$, we obtain that $\mathsf{M}$ is an equational pp expansion of $\mathsf{K}$. Then Theorem 12.4 implies that $\mathsf{M}$ is congruence preserving.
It remains to prove that $\mathsf{M}$ has the amalgamation property. Consider a pair of embeddings $f\colon\bm{A}\to\bm{B}$ and $g\colon\bm{A}\to\bm{C}$ with $\bm{A},\bm{B},\bm{C}\in\mathsf{M}$. As $\bm{B},\bm{C}\in\mathsf{M}=\mathbb{S}(\mathsf{K}[{L}_{\mathcal{F}}])$, there exist $\bm{B}^{\prime},\bm{C}^{\prime}\in\mathsf{K}[{L}_{\mathcal{F}}]$ and a pair of embeddings $i_{1}\colon\bm{B}\to\bm{B}^{\prime}$ and $i_{2}\colon\bm{C}\to\bm{C}^{\prime}$. Let $f^{\prime}\colon\bm{A}\to\bm{B}^{\prime}$ and $g^{\prime}\colon\bm{A}\to\bm{C}^{\prime}$ be the embeddings defined as $f^{\prime}=i_{1}\circ f$ and $g^{\prime}=i_{2}\circ g$. Observe that $f^{\prime}$ and $g^{\prime}$ can be viewed as embeddings $f^{\prime}\colon\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}}\to\bm{B}^{\prime}{\upharpoonright}_{{L}_{\mathsf{K}}}$ and $g^{\prime}\colon\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}}\to\bm{C}^{\prime}{\upharpoonright}_{{L}_{\mathsf{K}}}$. Furthermore, from $\bm{A},\bm{B}^{\prime},\bm{C}^{\prime}\in\mathsf{M}$ and Proposition 10.2 it follows that $\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}},\bm{B}^{\prime}{\upharpoonright}_{{L}_{\mathsf{K}}},\bm{C}^{\prime}{\upharpoonright}_{{L}_{\mathsf{K}}}\in\mathsf{K}$. Therefore, the assumption that $\mathsf{K}$ has the amalgamation property guarantees the existence of some $\bm{D}\in\mathsf{K}$ and embeddings $h_{f}\colon\bm{B}^{\prime}{\upharpoonright}_{{L}_{\mathsf{K}}}\to\bm{D}$ and $h_{g}\colon\bm{C}^{\prime}{\upharpoonright}_{{L}_{\mathsf{K}}}\to\bm{D}$ such that $h_{f}\circ f^{\prime}=h_{g}\circ g^{\prime}$.
@@
Recall that $\bm{D}\in\mathsf{K}$ and that $\mathsf{M}=\mathbb{S}(\mathsf{K}[{L}_{\mathcal{F}}])$ is a pp expansion of $\mathsf{K}$. Therefore, by Proposition 10.2 there exists $\bm{D}^{\prime}\in\mathsf{K}[{L}_{\mathcal{F}}]$ such that $\bm{D}\leqslant\bm{D}^{\prime}{\upharpoonright}_{{L}_{\mathsf{K}}}$. Let $h_{f}^{\prime}\colon\bm{B}^{\prime}{\upharpoonright}_{{L}_{\mathsf{K}}}\to\bm{D}^{\prime}{\upharpoonright}_{{L}_{\mathsf{K}}}$ and $h_{g}^{\prime}\colon\bm{C}^{\prime}{\upharpoonright}_{{L}_{\mathsf{K}}}\to\bm{D}^{\prime}{\upharpoonright}_{{L}_{\mathsf{K}}}$ be the embeddings obtained by composing $h_{f}$ and $h_{g}$, respectively, with the inclusion map of $\bm{D}$ into $\bm{D}^{\prime}{\upharpoonright}_{{L}_{\mathsf{K}}}$. Since $\bm{B}^{\prime},\bm{C}^{\prime},\bm{D}^{\prime}\in\mathsf{K}[{L}_{\mathcal{F}}]$, Proposition 9.5 implies that $h_{f}^{\prime}\colon\bm{B}^{\prime}\to\bm{D}^{\prime}$ and $h_{g}^{\prime}\colon\bm{C}^{\prime}\to\bm{D}^{\prime}$ are also embeddings. Moreover, from $h_{f}\circ f^{\prime}=h_{g}\circ g^{\prime}$ it follows that $h_{f}^{\prime}\circ f^{\prime}=h_{g}^{\prime}\circ g^{\prime}$. Together with $f^{\prime}=i_{1}\circ f$, and $g^{\prime}=i_{2}\circ g$, this yields
@@
To this end, consider $\bm{A}\in\mathsf{M}$ and observe that $\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}}\in\mathsf{K}$ by Proposition 10.2. From Proposition 2.10 it follows that
@@
It only remains to prove the last part of the statement. Suppose that $\mathsf{K}$ is a variety. From Theorem 10.3(ii) it follows that $\mathsf{M}$ is a quasivariety. Therefore, it remains to show that $\mathsf{M}$ is closed under homomorphic images. By Proposition 2.5 it suffices to prove that $\mathsf{Con}(\bm{A})=\mathsf{Con}_{\mathsf{M}}(\bm{A})$ for every $\bm{A}\in\mathsf{M}$. To this end, observe that for every $\bm{A}\in\mathsf{M}$,
@@
where the first step holds by Remark 12.2, the second follows from $\bm{A}{\upharpoonright}_{{L}_{\mathsf{K}}}\in\mathsf{K}$ (see Proposition 10.2) and the assumption that $\mathsf{K}$ is a variety, and the third from $\bm{A}\in\mathsf{M}$ and the assumption that $\mathsf{M}$ is a congruence preserving pp expansion of $\mathsf{K}$. Thus, $\mathsf{Con}(\bm{A})=\mathsf{Con}_{\mathsf{M}}(\bm{A})$, as desired. ∎